OpenStudy (anonymous):

X varies directly with the cube of w. The constant of proportionality is 4π/3. Relate the quantities.

OpenStudy (hunus):

\[x = k * w^{3}\] Where k = 4pi/3 \[x = \frac{4\pi*w^{3}}{3}\]

OpenStudy (anonymous):

You're awesome!